OpenStudy (wesdg1978):

f(x)=ax^3+bx^2+cx+d (a,b,c,d constant) and f(x)= (1/a)(x^2+(1/b)x+c) (a,b,c constant) How do you find f '(x)? The answer to the first one is: 12x(3x^2+1) but I don't know how to get there. Can someone help please?

OpenStudy (anonymous):

\[f(x)=ax^3+bx^2+cx+d\] Differentiate term by term, using the power rule for each term with an x, and the constant rule for d: \[f'(x)=3ax^{3-1}+2bx^{2-1}+cx^{1-1}+\frac{d}{dx}d\\ f'(x)=3ax^{2}+2bx^{1}+cx^{0}+0\\ f'(x)=3ax^{2}+2bx+c\] For the second function: \[f(x)=\frac{1}{a}\left(x^2+\frac{1}{b}x+c\right)\] Go ahead and distribute the coefficients, then differentiate term by term like in the first function: \[f(x)=\frac{1}{a}x^2+\frac{1}{ab}x+\frac{c}{a}\\ f'(x)=\frac{2}{a}x^{2-1}+\frac{1}{ab}x^{1-1}+\frac{d}{dx}\frac{c}{a}\\ f'(x)=\frac{2}{a}x^{1}+\frac{1}{ab}x^{0}+0\\ f'(x)=\frac{2}{a}x+\frac{1}{ab}\] I'm not sure how the answer for the first is supposed to be \(12x\left(3x^2+1\right)\), since you're not given any information about the constants \(a,b,c,d\). At first, I thought that maybe the two given functions are the same, but that's not possible, since \[ax^3+bx^2+cx+d\not= \frac{1}{a}\left(x^2+\frac{1}{b}x+c\right)\] because the left side is a cubic and the right side is a quadratic.